# Geometric Neural Network based on Phase Space for BCI decoding

### v1

@@ 5.1 Impact of augmentation on most and least responsive 5 subject of Cho2017
In the previous section, we explored the outcome obtained by SPDNet within the context of a reduced number of electrodes. Now, our attention shifts to a deeper exploration of this new algorithm, with a specific emphasis on the interpretability and computational performance of the method. To facilitate this comprehensive analysis, still in the context of three electrodes, we have chosen to concentrate on the examination of the five most and least responsive subjects within the Cho2017 dataset (the wider dataset in terms of the number of subjects), selected using the Minimum Distance to the Mean (MDM) algorithm [55] with covariance as a feature.
To analyze the performance, we use different pipelines, MDM [55], AugMDM [7], SPDNet, and SPDNet, each offering a distinct perspective on the impact of augmentation techniques on both traditional Machine Learning (ML) and advanced Deep Learning (DL) algorithms. Furthermore, we assess the performance using three different SPD features: covariance, instantaneous coherence, and imaginary coherence (Figure 3).
@@ 5.2 Interpretability
For analyzing the explainability of the SPDNet we use the GradCam++ algorithm [70], a technique designed for visualizing and explaining the decision-making process of neural networks. The GradCam++ output is essentially a visual representation that reveals which elements within the data are the most important for the model. GradCam++ has proven to be a robust method for spatial interpretability of EEG signals in a class-agnostic manner [71].
In our analysis, we focus on Subject 41 from the Cho2017 dataset, which is notable for being among the top 5 subjects that exhibit the most substantial performance improvement through the augmentation procedure. Similar findings also hold for other subjects. In particular, using covariance as a feature, we show a sample of the test dataset that was correctly classified by the SPDNet algorithm (left hand classified as left hand) but was misclassified by SPDNet 4 using the same feature. Similar findings have been observed for the FC estimators. This figure represents the output of the GradCam++ applied to the output of the ReEig layer.
In all three scenarios, the standard SPDNet places more emphasis on the diagonal terms of SPD matrices. On the other side, when employing SPDNet, a more comprehensive picture unfolds. While diagonal terms continue to play a pivotal role, what becomes increasingly evident is the significant contribution of the off-diagonal terms. In essence, this observation emphasizes that the SPDNet model leverages both individual electrode characteristics (diagonal terms) and the interplay between electrodes in time and space (off-diagonal terms) as crucial elements in its classification decision-making process.
To delve deeper into the interpretability of our proposed network, we employ a paired t-test to highlight statistically significant differences in the augmented covariance matrix (ACM) between left and right hands. Our analysis centers on subject 41 from Cho2017, specifically on the test dataset of the initial fold of the cross-validation, with consistent findings observed in the other folds. Given that 20 samples are available for each imagined movement, we are in the domain of application of the t-test.
Examining Figure 4 (d) and (e), it is evident that SPDNet places emphasis primarily on the statistically significant pair of electrodes, showing a clear parallel between the GradCam++ visualization and the t-test. In contrast, SPDNet not only considers significant interactions but also extends its focus to nearly significant interactions.
@@ 5.3 Convergence Behavior
Notably, our observations reveal that the loss function of SPDNet converges to zero at a notably accelerated pace and exhibits a considerably more stable trajectory compared to standard SPDNet. Moreover, while the loss function of SPDNet steadily approaches zero, the standard SPDNet experiences a more gradual descent, deviating from the path to complete convergence but ultimately plateauing around . These findings highlight that the capacity of SPDNet is limited and that the algorithm stops learning at a certain point.
Interestingly, when examining the ROC AUC graph, we observe that the validation scores of SPDNet converge consistently to higher values. Notably, this convergence is achieved even faster compared to the SPDNet standard, requiring only of the iterations. This implies that SPDNet attains a comparable performance to SPDNet but with a significantly reduced number of iterations, making it an efficient and effective choice for real-life applications.
@@ 5.4 Computational Performances
Interestingly our analysis shows that the SPDNetexhibits the longer computational time among the evaluated algorithms while simultaneously standing out as one of the most environmentally friendly in terms of its impact (see Fig 5.4). This intriguing juxtaposition can be attributed to the fact that, during each fold of the cross-validation process, there’s a necessity to optimize the model’s hyper-parameters, which naturally augments the computational time. However, the MDOP algorithm consumes fewer resources than gradient descent procedures, thus resulting in a comparatively smaller environmental impact. In fact, the use of augmented procedure introduces the complication of two hyper-parameters but nevertheless manages to keep the number of trainable parameters of the model contained (see Table 5.4). It’s plausible that our algorithm has yet to reach its optimal performance. Further investigation is needed to improve the optimization strategies.
@@
Avg. Time (s)
Emission (g )
DeepNet
21.77
0.39
ShallowNet
15.10
0.25
EEGNet-8,2
14.96
0.25
EEGITNet
28.44
0.48
EEGTCNet
28.83
0.47
EEGNeX
28.15
0.46
SPDNet
6.80
0.07
SPDNet
30.16
0.22
@@
tableAverage computational time and equivalent consumption over the best five subjects of Cho2017
Stage
Layer
Output
Parameters
Input
[1, 3, 1537]
0
Augmentation
AugmentedDataset
[1, 18, 1393]
0
Covariance
Covariances
[1, 18, 18]
0
SPDNet
BiMap
[1, 9, 9]
162
ReEig
[1, 9, 9]
0
LogEig
[1, 45]
0
Classification
Linear
[1, 2]
45
Total
207
[…]
Here, it is reported for the specific case of order .
@@ 5.5 Limitation and Future Directions
The existing SPDNet, as employed thus far, utilizes MDOP for hyper-parameter selection, offering computational efficiency compared to grid-search methods [7]. However, the solutions obtained with MDOP were sub-optimal compared with using a more extensive hyper-parameter selection. A potential avenue for future exploration involves streamlining and accelerating the grid-search process to harness the capabilities of the SPDNet architecture fully.
@@
An intriguing direction arises from the insights gained in the explainability study, revealing that SPDNetpredominantly focuses on diagonal elements. This suggests the possibility of employing a Region of Interest approach, where the network initially prioritizes the main diagonals while treating the remainder of the matrix as background.
@@ 6 Conclusion
Validation of our approach involves leveraging nearly 100 subjects from several open datasets, a task facilitated by the MOABB framework. The resulting SPDNetalgorithm outperforms existing DL algorithms in BCI-EEG classification and offers some explainability elements through GradCam++ visualization. Remarkably, this algorithm requires a modest number of trainable parameters and exhibits a lower environmental impact, measured by equivalent emissions.
@@ Appendix A Results using Coherence as feature
In this section, we shift our focus to report results due to the use of Imaginary and Instantaneous coherence features as input. Specifically, we aim to draw a comparison between the SPDNetresults and the non DL state-of-the-art pipelines (usually created for covariance but here applied on coherence). It’s worth noting that, in the subsequent analysis, we abstain from comparing these results with the performance of state-of-the-art DL models, given that the employment of coherence-based features typically yields lower results across the board. Nevertheless, we find it particularly intriguing to present the outcomes achieved by SPDNetin the context of coherence-based features since we observe that SPDNetdemonstrates a statistically significant improvement. This shows the potential and significance of SPDNetin the broader landscape of signal processing and pattern recognition.
@@
In this study the situation is even rosier: SPDNet, used with either imaginary or instantaneous coherence features, emerges as best pipeline across all datasets. When utilizing imaginary coherence, SPDNet alone already achieves statistically superior results compared to state-of-the-art approaches employing the same features. However, the augmentation procedure introduces a significant performance boost, a finding supported by the meta-effect analysis shown in Fig A1 (c).
In the case of instantaneous coherence, on the other hand, the impact of the augmentation procedure is even more significant. In fact, with this feature, the standard SPDNet fails to exhibit statistically superior performance in comparison to the state-of-the-art pipelines. SPDNet, on the other hand, stands out as the top performing pipeline, showing that the augmentation procedure bring a statistically significant influence on classification outcomes (see Fig A2 (a)).

### v2

@@ 5.1 Impact of augmentation on most and least responsive 5 subject of Cho2017
In the previous section, we explored the outcome obtained by Phase-SPDNet within the context of a reduced number of electrodes. Now, our attention shifts to a deeper exploration of this new algorithm, with a specific emphasis on the interpretability and computational performance of the method. To facilitate this comprehensive analysis, still in the context of three electrodes, we have chosen to concentrate on the examination of the five most and least responsive subjects within the Cho2017 dataset (the wider dataset in terms of the number of subjects), selected using the Minimum Distance to the Mean (MDM) algorithm [63] with covariance as a feature.
To analyze the performance, we use different pipelines, MDM [63], ACM+MDM [7], SPDNet, and Phase-SPDNet , each offering a distinct perspective on the impact of augmentation techniques on both traditional ML and advanced DL algorithms. Furthermore, we assess the performance using three different SPD features: covariance, instantaneous coherence, and imaginary coherence (Figure 4).
@@ 5.2 Interpretability
For analyzing the explainability of the Phase-SPDNet we use the GradCam++ algorithm [86], a technique designed for visualizing and explaining the decision-making process of neural networks. The GradCam++ output is essentially a visual representation that reveals which elements within the data are the most important for the model. GradCam++ has proven to be a robust method for spatial interpretability of EEG signals in a class-agnostic manner [87].
In our analysis, we focus on Subject 41 from the Cho2017 dataset, which is notable for being among the top 5 subjects that exhibit the most substantial performance improvement through the augmentation procedure. Similar findings also hold for other subjects. In particular, using covariance as a feature, we show a sample of the test dataset that was correctly classified by the Phase-SPDNet algorithm (left hand classified as left hand) but was misclassified by SPDNet using the same feature, as shown in Figure 5. Similar findings have been observed for the FC estimators. This figure represents the output of the GradCam++ applied to the output of the ReEig layer.
In all three scenarios, the standard SPDNet places more emphasis on the diagonal terms of SPD matrices. On the other side, when employing Phase-SPDNet , a more comprehensive picture unfolds. While diagonal terms continue to play a pivotal role, what becomes increasingly evident is the significant contribution of the off-diagonal terms. In essence, this observation emphasizes that the Phase-SPDNet model leverages both individual electrode characteristics (diagonal terms) and the interplay between electrodes in time and space (off-diagonal terms) as crucial elements in its classification decision-making process.
To investigate deeper into the interpretability of our proposed network, we employ a paired t-test to highlight statistically significant differences in the augmented covariance matrix (ACM) between left and right hands. Our analysis centers on subject 41 from Cho2017, specifically on the test dataset of the initial fold of the cross-validation, with consistent findings observed in the other folds. Given that 20 samples are available for each imagined movement, we are in the domain of application of the t-test.
Examining Figure 5 (d) and (e), it is evident that SPDNet places emphasis primarily on the statistically significant pair of electrodes, showing a clear parallel between the GradCam++ visualization and the t-test. In contrast, Phase-SPDNet not only considers significant interactions but also extends its focus to nearly significant interactions.
@@ 5.3 Convergence Behavior
Notably, our observations reveal that the loss function of Phase-SPDNet converges to zero at a notably accelerated pace and exhibits a considerably more stable trajectory compared to standard SPDNet. Moreover, while the loss function of Phase-SPDNet steadily approaches zero, the standard SPDNet experiences a more gradual descent, deviating from the path to complete convergence but ultimately plateauing around . These findings highlight that the capacity of SPDNet is limited and that the algorithm stops learning at a certain point.
Interestingly, when examining the ROC AUC graph, we observe that the validation scores of Phase-SPDNet converge consistently to higher values. Notably, this convergence is achieved even faster compared to the SPDNet standard, requiring only of the iterations. This implies that Phase-SPDNet attains a comparable performance to SPDNet but with a significantly reduced number of iterations, making it an efficient and effective choice for real-life applications.
@@ 5.4 Computational Performances
Interestingly, our analysis shows that the Phase-SPDNet exhibits the longer computational time among the evaluated algorithms while simultaneously standing out as one of the most environmentally friendly in terms of its impact (see Fig 5.4). This intriguing juxtaposition can be attributed to the fact that, during each fold of the cross-validation process, there’s a necessity to optimize the model’s hyper-parameters, which naturally augments the computational time. However, the MDOP algorithm consumes fewer resources than gradient descent procedures, thus resulting in a comparatively smaller environmental impact. In fact, the use of augmented procedure introduces the complication of two hyper-parameters but nevertheless manages to keep the number of trainable parameters of the model contained (see Table 5.4). It’s plausible that our algorithm has yet to reach its optimal performance. Further investigation is needed to improve the optimization strategies.
@@
| Avg. Time (s) | Emission (g ) | |
| DeepNet | 21.77 | 0.39 |
| ShallowNet | 15.10 | 0.25 |
| EEGNet-8,2 | 14.96 | 0.25 |
| EEGITNet | 28.44 | 0.48 |
| EEGTCNet | 28.83 | 0.47 |
| EEGNeX | 28.15 | 0.46 |
| DynSpat+ShallowNet | 34.42 | 0.41 |
| DynSpat+EEGNet | 34.09 | 0.45 |
| SPDNet | 6.80 | 0.07 |
| Phase-SPDNet | 30.16 | 0.22 |
@@ 5.5 Limitation and Future Directions
The existing Phase-SPDNet , as employed thus far, utilizes MDOP for hyper-parameter selection, offering computational efficiency compared to grid-search methods [7]. However, the solutions obtained with MDOP were sub-optimal compared with using a more extensive hyper-parameter selection. A potential avenue for future exploration involves streamlining and accelerating the grid-search process to harness the capabilities of the Phase-SPDNet architecture fully.
@@
An intriguing direction arises from the insights gained in the explainability study, revealing that Phase-SPDNet predominantly focuses on diagonal elements. This suggests the possibility of employing a Region of Interest approach, where the network initially prioritizes the main diagonals while treating the remainder of the matrix as background.
@@ 6 Conclusion
Validation of our approach involves leveraging nearly 100 subjects from several open datasets, a task facilitated by the MOABB framework. The resulting Phase-SPDNet algorithm outperforms existing DL algorithms in BCI-EEG classification and offers some explainability elements through GradCam++ visualization. Remarkably, this algorithm requires a modest number of trainable parameters and exhibits a lower environmental impact, measured by equivalent emissions.
@@ Appendix A Results using Coherence as feature
In this section, we shift our focus to reporting results due to the use of Imaginary and Instantaneous coherence features as input. Specifically, we aim to draw a comparison between the Phase-SPDNet results and the non DL state-of-the-art pipelines (usually created for covariance but here applied on coherence). It’s worth noting that, in the subsequent analysis, we abstain from comparing these results with the performance of state-of-the-art DL models, given that the employment of coherence-based features typically yields lower results across the board. Nevertheless, we find it particularly intriguing to present the outcomes achieved by Phase-SPDNet in the context of coherence-based features since we observe that Phase-SPDNet demonstrates a statistically significant improvement. This shows the potential and significance of Phase-SPDNet in the broader landscape of signal processing and pattern recognition.
@@
In this study, the situation is even rosier: Phase-SPDNet , used with either imaginary or instantaneous coherence features, emerges as the best pipeline across all datasets. When utilizing imaginary coherence, SPDNet alone already achieves statistically superior results compared to state-of-the-art approaches employing the same features. However, the augmentation procedure introduces a significant performance boost, a finding supported by the meta-effect analysis shown in Fig A1 (c).
In the case of instantaneous coherence, on the other hand, the impact of the augmentation procedure is even more significant. In fact, with this feature, the standard SPDNet fails to exhibit statistically superior performance in comparison to the state-of-the-art pipelines. Phase-SPDNet , on the other hand, stands out as the top-performing pipeline, showing that the augmentation procedure brings a statistically significant influence on classification outcomes (see Fig A2 (a)).